OpenStudy (anonymous):

Choose the procedure you would use to solve the equation 5x + 3 = 28 for x.

OpenStudy (cwrw238):

subtract 3 from both sides of equation then divide both sides by 5 to get value of x

OpenStudy (anonymous):

5x+3=28 =>5x=28-3 =>5x=25 =>x=25/5 =>x=5 hope this helps